# Personalizing Agent Privacy Decisions via Logical Entailment

### v2

@@ 3.2. Baseline Approaches
Using an LLM to address the privacy decision problem defined above requires addressing a key challenge: how to design the prompt for the LLM. Here, we discuss two baseline approaches to prompt the LLMs to make privacy judgments along with their limitations.
Norm-based Baseline:
The first choice for prompting the LLM is to utilize Privacy Norms. In this approach, the prompt contains three parts. The first part is the instruction to the LLM to make privacy decisions, the second part is the incoming request, and the third part includes the privacy norms. These privacy norms, akin to CI-inspired approaches, can include the general population’s overall acceptability on particular action (fan2024goldcoin). For instance, determining if a user should share a partial SSN with a bank would rely on broad societal acceptability. However, such an approach is limited because general norms lack the granularity required to capture specific, personalized user privacy decisions.
Privacy Preferences-based Baseline:
To address this limitation, an alternative approach is to prompt the LLM explicitly with the user’s privacy preferences. In this approach, the prompt also includes three parts: the general task instruction, the incoming request, and the user’s privacy preferences. These preferences can be based on rules or policies that the user specify, similar to groschupp2025can, or they can include a list of the user’s privacy judgments on prior requests (e.g., leveraging a past decision about sharing a full SSN to infer judgment of incoming request).
@@
Appropriateness judgments generated directly by LLMs remain unreliable due to limited logical reasoning capabilities (lin2025zebralogic) and the potential for hallucinations that deviate from system prompts (ji2023survey; kalai2025language; zhang2023siren). Consequently, these judgments are not guaranteed to align with user privacy preferences, resulting in erroneous judgments that limit effective personalization.
@@
Finally, because these methods do not align with user preferences and are not inherently interpretable, they hinder user agency. While personalized agents should aim to mitigate privacy fatigue, this objective must not come at the cost of compromising user awareness of and control over how decisions are made on their behalf (mitchell2025fully).
@@ 3.3. Requirements
    Alignment. The privacy decisions made by the agent on behalf of the users need to accurately reflect the user’s prior privacy decisions. Alignment is crucial for effective personalization of agents.
@@
    Interpretability. The agent needs to provide a reasoning that is transparent and traceable so that the user stays in the loop throughout the entire privacy-decision process. Interpretability maintains user trust in the system.
@@
    User Agency. In scenarios where the agent lacks enough information from the user’s prior privacy decisions to make a judgment, the agent must determine when it should escalate the request to the user for assistance. Importantly, the agent needs to accomplish the right tradeoff of automated privacy decision vs user control.
@@ 4. ARIEL: Privacy Judgments with Entailment
We propose ARIEL, a system to enable more reliable and structured reasoning for personalized privacy judgments. Below, we describe ARIEL in more detail.
Figure 3. High-level overview of ARIEL, which can be broken down into three components. ① ARIEL generates an ontology for each user based on their prior privacy judgments from the knowledge base using the LLM. ② Once ARIEL receives an incoming request, it goes through each prior request in the user’s knowledge base to determine if entailment holds. It uses the LLM and the generated ontologies to map the parameters in both requests to correspond levels in the ontologies. ③ A set of rules are applied to determine if entailment holds between each mapped prior requests and the mapped prior request. If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
@@ 4.1.1. Entailment
This approach addresses the limitations of LLM-based baselines, which do not provide a logical way to infer user judgments from prior decisions. In contrast, entailment treats prior judgments as logical premises that constrain future decisions, offering an interpretable and controllable method to making privacy decisions. This example highlights how entailment enables privacy decision based on prior requests.
@@ 4.1.2. Design
    Ontology Generation: use an LLM to generate relevant ontologies for each user over the parameters on incoming and prior requests.
@@
The three components of ARIEL operate in two phases. In the offline phase, Ontology Generation creates ontologies for each user based on the user’s prior privacy judgments and general privacy norms encoded in the LLM. The online phase processes each incoming request: first, Ontology Mapping selects relevant prior requests with user judgments from the user’s knowledge base, then maps the parameters in each request to the generated ontologies. This step provides a unified representation of the parameters in both the incoming and prior requests, enabling the entailment step. Next, Entailment Logic applies a set of rules to determine if entailment holds between each mapped prior requests and the mapped incoming request. If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
@@ 4.2. Ontology Generation
ARIEL instructs the LLM to generate an ontology for each request parameter, excluding the data subject, which is always the same for each request. The LLM generates the ontology based on a hierarchical relationship. We detail each ontology with its corresponding relationship below (please refer to Appendix A.5 for the exact ontology generation prompts):
@@
    Data Recipient. Represents the user’s trust in the recipient and the recipient’s authority. Trust and authority decrease at higher levels (Level 1 has highest trust/authority).
@@
    Transmission Principle. Represents the transmission purpose and safeguards. Safeguard strength and purpose relevance decrease at higher levels (Level 1 has strongest safeguards/relevance).
Each hierarchical relationship is derived from two sources: (1) User’s prior privacy judgments: Relationships are inferred directly from the user’s prior privacy judgments ($D_{u}$). For instance, if a user approves a request involving parents but denies an identical request involving friends, ARIEL infers a higher trust level for parents. This ensures the ontology is personalized. (2) General norms: For relationships not present in $D_{u}$, ARIEL leverages the LLM’s encoded societal norms. For example, the LLM can infer that a typical user generally trusts family more than acquaintances. Figure 4 illustrates a data type ontology generated by Gemini 2.5 Pro using this method.
@@ 4.3. Ontology Mapping
In the online phase, ARIEL evaluates an incoming request $R_{u}$ by comparing it against prior requests. The challenge is that the privacy contexts can diverge significantly if requests differ by multiple parameters simultaneously, making direct entailment unreliable. For example, sharing full SSN with a bank to open a checking account is an entirely different context than sharing full SSN with an IRS agent to file taxes. Furthermore, even when requests are similar, the agent requires a structured process for comparing differing parameter values based on some hierarchal relationship in order to determine the entailment (e.g., comparing data recipients of parents vs friends).
@@
Neighboring Requests, To ensure that comparison is done on contextually similar requests, we first select prior requests that ’neighbor’ the incoming request, denoting the neighboring set as $E$. More precisely, $E$ contains prior requests $R^{i}_{u}$ that have a Hamming distance $d$ of at most one with the incoming request $R_{u}$. This is equivalent to ensuring that both requests differ by only one parameter (line 6 in Algorithm 1).
Ontology Mapping. To rigorously compare the differing parameters, we introduce an operator $\leq_{O}$ to formalize the relationship between the parameters of two distinct requests, $R$ and $R^{\prime}$, based on their mapped levels within $O$. Specifically, $R^{\prime}\leq_{O}R$ holds if every parameter in $R^{\prime}$ is ontologically ’less than or equal to’ the corresponding parameter in $R$. We formally define this as:
@@
Let $R=\{d_{u},u,s,r,t\}$, $R^{\prime}=\{d^{\prime}_{u},u,s^{\prime},r^{\prime},t^{\prime}\}$ be two requests defined in Definition 3.2 that are partially ordered by the set of ontologies $O$. Then $R^{\prime}\leq_{O}R$ is true if $\forall p\in R,\forall p^{\prime}\in R^{\prime}$ either $p^{\prime}<_{o}p$ or $p^{\prime}=_{o}p$ for the relevant ontology $o$.
Then ARIEL iterates through each neighboring prior request $R^{\prime}_{u}$. For each pair, the agent retrieves the differing parameter values, $p$ (from $R$) and $p^{\prime}$ (from $R^{\prime}_{u}$). These values are then mapped to the relevant ontology $o\in O$ (line 12 in Algorithm 1). For this ontological mapping, we use the LLM $\theta$ to determine the relevant ontology $o$ and map $p$ and $p^{\prime}$ to their corresponding levels in $o$
@@ 4.4. Entailment Logic
Let $R,R^{\prime}$ be two requests with $l^{\prime}$ being an appropriateness label for $R^{\prime}$. And let $O$ be a set of ontologies. We say that the entailment $(R^{\prime},l^{\prime})\implies(R,l^{\prime})$ holds if one of the following two cases hold:
@@
ARIEL checks if entailment holds using Definition 4.3 (line 13 in Algorithm 1). We note that multiple neighboring requests can be entailed, with potentially opposing judgments. This could be due to inconsistency in the user’s prior judgments or incomplete information from the ontology. To address this, the agent goes through each neighboring request and keeps a running vote of the appropriate and inappropriate judgments that were entailed. Whichever appropriateness receives the majority vote is the final judgment. If there is no majority, then the agent outputs ’undetermined’ (line 19 in Algorithm 1), signifying that the incoming request should be escalated to the user for further assistance.
Lastly, we note that in scenarios where multiple appropriate and inappropriate prior requests are entailed, ARIEL can be modified for the agent to escalate the incoming request to the user instead of employing majority vote.
@@ 5.1. Datasets for Data-sharing Requests
We use existing datasets that elicit privacy preferences through factorial vignette surveys based on CI. Participants are presented with data-sharing scenarios and asked to indicate their acceptance of each one (abdi2021privacy; apthorpe2018discovering; apthorpe2019evaluating; bhatia2018empirical; hoyle2020privacy; martin2016measuring; shvartzshnaider2016learning). To simulate a user with prior privacy judgments, each of the data-sharing scenarios are converted into requests with appropriateness judgments from the participants. Then, we divided a participant’s set of answered requests into two pairwise disjoint subsets. The first subset was treated as observed prior requests with the user’s answers as the judgments. The second subset represents incoming requests where the participant’s answers were held out as the ground truth for evaluation. Our evaluation was able to employ two such datasets, which we repurposed for our setup as detailed below. See Figure 5 for example requests from both datasets.
@@ 5.1.1. Smart Home Personal Assistant (SPA) dataset
The SPA dataset [^1] Dataset can be accessed here: https://osf.io/63wsm/overview (abdi2021privacy) consists of factorial vignette surveys covering different data-sharing scenarios to understand privacy norms in the SPA ecosystem. The scenarios typically involve a voice assistant provider sending user data to individuals/entities who have access to the user’s SPA ecosystem, under a certain purpose and condition. The data subject of a request is always ’user’ and the data sender is always ’assistant provider’. The total number of participants is 1,730, with each participant randomly assigned to answer a subset of the questions. This resulted in approximately 180 different scenarios answered by each participant.
Due to the relatively large number of participant responses from the SPA dataset, we restrict our evaluation to a set of randomly sampled 500 users. To ensure a consistent experimental setup across all users, we standardized the response size by randomly selecting 70 responses per user. For each of the 500 users, we partitioned the responses into 60 prior requests and 10 incoming requests. This yields a total of 5,000 incoming requests, consisting of 1,765 appropriate and 3,235 inappropriate labels. We also perform an ablation study on varying the number of prior requests in Section 6.2.
Lastly, in the dataset, the participant responses are in 1-5 Likert scale where 1-2 correspond to strongly and somewhat unacceptable, 3 corresponds to neutral, and 4-5 correspond to somewhat and strongly acceptable. We convert the responses with 1-2 scores to ”Inappropriate” and the responses with 4-5 to ”Appropriate”. We filter the remaining responses with scores of 3.
@@
The participant responses in the dataset contain numerical values, where 1 corresponds to Yes, 2 corresponds to No, and 3-5 correspond to Does not Make Sense options. Specifically, 3 is ’The sender is unlikely to have the information, 2 is ’The receiver would already have the information, and 3 is ’the question is ambiguous.’ We convert responses with 1 to ”Appropriate” and with 2 to ”Inappropriate”. We filter out the remaining responses with 3-5 scores. Lastly, we filter out users that have less than 70 appropriateness judgments, so each user has 60 prior requests and 10 incoming requests. In total, there are 302 users and 30210=3020302*10=3020 total incoming requests, consisting of 1378 appropriate labels and 1642 inappropriate labels.
@@ 5.2. Baselines
5.2.1. Norms-based Baseline
Privacy norms capture the general population’s accepted appropriateness of information flows. There are two prompts that we consider.
@@ 5.2.1. Norms-based Baseline
    zero-shot. We construct the prompt to utilize the out-of-the-box privacy norm awareness of LLMs, following from recent prior work on evaluating CI-compliance in LLMs (mireshghallah2024can; shao2024privacylens). The prompt directs the LLM to judge the appropriateness of each information flow based on its knowledge of general privacy norms. Prompts can be found in Appendix A.1.
@@
    privacy norms. We construct the prompt such that the LLM’s privacy judgment is now based on a provided subset of representative privacy norms. These extracted privacy norms are taken directly from the analysis provided in the papers of each dataset. Table 2 contains a subset of the general privacy from both evaluation datasets that were used by the LLMs. The prompt directs the LLM to judge the appropriateness of each information flow based on the provided general privacy norms. Prompts can be found in A.2.

### v3

@@ 3.2. Baseline Approaches
Using an LLM to address the privacy decision problem defined above requires addressing a key challenge: how to design the prompt for the LLM. Here, we discuss two baseline approaches to prompt the LLM to make privacy judgments, along with their limitations.
Privacy Norms-based Baseline.
The first choice for prompting the LLM is to utilize Privacy Norms. In this approach, the prompt contains three parts. The first part is the instruction to the LLM to make privacy decisions, the second part includes the privacy norms, and the third part is the incoming request. These privacy norms, akin to CI-inspired approaches, can include the general population’s overall acceptability of a particular action (fan2024goldcoin). For instance, determining whether a user should share a partial SSN with a bank would rely on broad societal acceptability. However, such an approach is limited because general norms lack the granularity required to capture user-specific, personalized privacy decisions.
Privacy Preferences-based Baseline.
To address this limitation, an alternative approach is to prompt the LLM explicitly with the user’s privacy preferences. In this approach, the prompt also includes three parts: the general task instruction, the user’s privacy preferences, and the incoming request. These preferences can be based on rules or policies that the user specifies, similar to groschupp2025can, or they can include a list of the user’s privacy judgments on prior requests (e.g., leveraging a past decision about sharing a full SSN to infer the judgment of incoming request) (wijesekera2017feasibility; colnago2022concern).
@@
Appropriateness judgments generated directly by LLMs remain unreliable due to limited logical reasoning capabilities (lin2025zebralogic) and the potential for hallucinations that deviate from system prompts (ji2023survey; kalai2025language; zhang2023siren). Consequently, these judgments are not guaranteed to align with user privacy preferences, resulting in errors that limit effective personalization.
@@
Finally, because these methods do not align with user preferences and are not inherently interpretable, they hinder user agency. While personalized agents should aim to mitigate privacy fatigue, this objective must not come at the cost of compromising user awareness of and control over how decisions are made on their behalf (mitchell2025fully; zhang2025autonomy).
@@ 3.3. Requirements
    Alignment. The privacy decisions made by the agent on behalf of the user need to accurately reflect the user’s prior privacy decisions. Alignment is crucial for effective personalization of agents.
@@
    Interpretability. The agent needs to provide reasoning that is transparent and traceable so that the user stays in the loop throughout the privacy-decision process. Interpretability maintains user trust in the system.
@@
    User Agency. In scenarios where the agent lacks enough information from the user’s prior privacy decisions to make a judgment, the agent must determine when it should escalate the request to the user for assistance. Importantly, the agent needs to strike the right trade-off between automated privacy decisions and user control.
@@ 4. ARIEL: Privacy Judgments with Entailment
We propose ARIEL, a framework enabling aligned, structured, and interpretable reasoning for personalized privacy judgments. ARIEL operates as a specialized module for privacy decisions within an end-to-end agentic system (see Figure 1). We describe the framework in detail below.
Figure 3. High-level overview of ARIEL, which can be broken down into three components. ① ARIEL uses an LLM to generate an ontology for each user based on the user’s prior privacy judgments on data-sharing requests. ② Once ARIEL receives an incoming request, it goes through each prior request from memory to determine if entailment holds. The LLM maps the parameters in both requests to corresponding levels in the generated ontologies. ③ A set of rules is applied to determine whether entailment holds between each mapped prior request and the mapped incoming request. If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
@@ 4.1.1. Entailment
This approach addresses the limitations of LLM-based baselines, which do not provide a logical way to infer user judgments from prior decisions. In contrast, entailment treats prior judgments as logical premises that constrain future decisions, offering an interpretable and controllable method for making privacy decisions. The following example highlights how entailment enables privacy decisions based on prior requests.
@@ 4.1.2. Design
    Ontology Generation: use an LLM to generate relevant ontologies for each user over the parameters in each prior request and incoming request.
@@
The three components of ARIEL operate in two phases. In the offline phase, Ontology Generation creates ontologies for each user based on the user’s prior privacy judgments and the LLM’s societal norm awareness. The online phase processes each incoming request: first, Ontology Mapping selects relevant prior requests with user judgments from memory, then maps the parameters in each request to the generated ontologies. This step provides a unified representation of the parameters in both the incoming and prior requests, enabling the entailment step. Next, Entailment Logic applies a set of rules to determine if entailment holds between each mapped prior request and the mapped incoming request. If the incoming request is not entailed by any prior requests, ARIEL escalates the incoming request to the user.
@@ 4.2. Ontology Generation
ARIEL instructs the LLM to generate an ontology for each request parameter, excluding the data subject, which is always the same for each request. The LLM generates each ontology based on a hierarchical relationship. We detail each ontology with its corresponding relationship below (please refer to Appendix A.5 for the exact ontology generation prompts):
@@
    Data Recipient. Represents the user’s trust in and the authority of the recipient. Trust and authority decrease at higher levels (Level 1 has highest trust/authority).
@@
    Transmission Principle. Represents the transmission purpose and safeguards. Safeguard strength and purpose relevance decrease at higher levels (Level 1 has the strongest safeguards/relevance).
Each hierarchical relationship is derived from two sources: (1) User’s prior privacy judgments: Relationships are inferred directly from the user’s prior privacy judgments ($D_{u}$). For instance, if a user approves a request involving parents but denies an identical request involving friends, ARIEL infers a higher trust level for parents. This ensures the ontology is personalized. (2) General norms: For relationships not present in $D_{u}$, ARIEL leverages the LLM’s societal norm awareness. For example, the LLM can infer that a typical user generally trusts family more than acquaintances. Figure 4 illustrates a data type ontology generated by Gemini 2.5 Pro for a particular user in the evaluation dataset using this method.
@@ 4.3. Ontology Mapping
In the online phase, ARIEL evaluates an incoming request $R_{u}$ by comparing it against prior requests. The challenge is that the privacy contexts can diverge significantly if requests differ by multiple parameters simultaneously, making direct entailment unreliable. For example, sharing full SSN with a bank to open a checking account is an entirely different context than sharing full SSN with an IRS agent to file taxes. Furthermore, even when requests are similar, the agent requires a structured process for comparing differing parameter values based on some hierarchical relationship in order to determine the entailment (e.g., comparing data recipients of parents vs friends).
@@
Neighboring Requests. To ensure that comparisons are made between contextually similar requests, we first select prior requests that ’neighbor’ the incoming request, denoting the neighboring set as $E$. More precisely, $E$ contains prior requests $R^{i}_{u}$ whose Hamming distance $d$ from the incoming request $R_{u}$ is at most one. This is equivalent to ensuring that both requests differ by only one parameter (line 6 in Algorithm 1).
Ontology Mapping. To rigorously compare the differing parameters, we introduce an operator $\leq_{O}$ to formalize the relationship between the parameters of two distinct requests, $R$ and $R^{\prime}$, based on their mapped levels within $O$. Specifically, $R^{\prime}\leq_{O}R$ holds if every parameter in $R^{\prime}$ is ontologically ”less than or equal to” the corresponding parameter in $R$. We formally define this as:
@@
Let $R=\{d_{u},u,s,r,t\}$, $R^{\prime}=\{d^{\prime}_{u},u,s^{\prime},r^{\prime},t^{\prime}\}$ be two requests defined in Definition 3.2 that are partially ordered by the set of ontologies $O$. Then $R^{\prime}\leq_{O}R$ holds if, for each corresponding parameter $p\in R$ and $p^{\prime}\in R^{\prime}$, either $p^{\prime}<_{o}p$ or $p^{\prime}=_{o}p$ under the relevant ontology $o\in O$.
Then ARIEL iterates through each neighboring prior request $R^{\prime}_{u}$. For each pair, the agent retrieves the differing parameter values, $p$ (from $R_{u}$) and $p^{\prime}$ (from $R^{\prime}_{u}$). These values are then mapped to the relevant ontology $o\in O$ (line 12 in Algorithm 1). For this ontological mapping, we use the LLM $\theta$ to determine the relevant ontology $o$ and map $p$ and $p^{\prime}$ to their corresponding levels in $o$.
@@ 4.4. Entailment Logic
Let $R,R^{\prime}$ be two requests with $l^{\prime}$ being an appropriateness label for $R^{\prime}$. Let $O$ be a set of ontologies. We say that the entailment $(R^{\prime},l^{\prime})\implies(R,l^{\prime})$ holds if one of the following two cases holds:
@@
ARIEL checks if entailment holds using Definition 4.3 (line 13 in Algorithm 1). We note that multiple neighboring requests can be entailed, with potentially opposing judgments. This could be due to inconsistency in the user’s prior judgments or incomplete information from the ontology. To address this, the agent goes through each neighboring request and keeps a running vote of the appropriate and inappropriate judgments that were entailed. Whichever appropriateness label receives the majority vote is the final judgment. If there is no majority, the agent outputs ’undetermined’ (line 19 in Algorithm 1), signifying that the incoming request should be escalated to the user for further assistance.
Lastly, we note that in scenarios where multiple appropriate and inappropriate prior requests are entailed, ARIEL can be modified so that the agent to escalate the incoming request to the user instead of employing a majority vote.
@@ 5.1. Datasets for Data-sharing Requests
We use existing datasets that elicit privacy preferences through factorial vignette surveys based on CI. Participants are presented with data-sharing scenarios and asked to indicate their acceptance of each one (abdi2021privacy; apthorpe2018discovering; apthorpe2019evaluating; bhatia2018empirical; hoyle2020privacy; martin2016measuring; shvartzshnaider2016learning). To simulate a user with prior privacy judgments, each data-sharing scenario is converted into a request (Definition 3.2) with appropriateness judgments directly received from the participants. Then, we divided a participant’s set of answered requests into two pairwise disjoint subsets. The first subset was treated as observed prior requests with the user’s answers as the judgments. The second subset represents incoming requests where the participant’s answers were held out as the ground truth for evaluation. Our evaluation employs two such datasets, which we repurposed for our setup as detailed below. See Figure 5 for example requests from both datasets.
@@ 5.1.1. Smart Home Personal Assistant (SPA) dataset
The SPA dataset [^1] Dataset can be accessed here: https://osf.io/63wsm/overview (abdi2021privacy) consists of factorial vignette surveys covering different data-sharing scenarios to understand privacy norms in the SPA ecosystem. The scenarios typically involve a voice assistant provider sending user data to individuals/entities who have access to the user’s SPA ecosystem, under a certain purpose and condition. The data subject of a request is always ”user” and the data sender is always ”assistant provider”. The total number of participants is 1,730, with each participant randomly assigned to answer a subset of the questions. This resulted in approximately 180 different scenarios answered by each participant.
Due to the relatively large number of participant responses from the SPA dataset, we restrict our evaluation to a set of randomly sampled 500 users. To ensure a consistent experimental setup across all users, we standardized the response size by randomly selecting 70 responses per user. For each of the 500 users, we partitioned the responses into 60 prior requests and 10 incoming requests. This yields a total of 50010=5000500*10=5000 incoming requests, consisting of 1,765 appropriate and 3,235 inappropriate labels. We also perform an ablation study on varying the number of prior requests in Section 6.2.
Lastly, in the dataset, the participant responses are on a 1-5 Likert scale where 1 and 2 correspond to strongly and somewhat unacceptable responses, respectively, 3 corresponds to neutral, and 4 and 5 correspond to somewhat and strongly acceptable responses, respectively. We convert the responses with 1-2 scores to ”Inappropriate” and the responses with 4-5 to ”Appropriate”. We filter the remaining responses with scores of 3.
@@
The participant responses in the dataset contain numerical values, where 1 corresponds to Yes, 2 corresponds to No, and 3-5 correspond to Does not Make Sense options. Specifically, 3 is ”The sender is unlikely to have the information”, 4 is ”The receiver would already have the information”, and 5 is ”the question is ambiguous.” We convert responses with 1 to ”Appropriate” and with 2 to ”Inappropriate”. We filter out the remaining responses with 3-5 scores. Lastly, we filter out users who have fewer than 70 appropriateness judgments, so each user has 60 prior requests and 10 incoming requests. In total, there are 302 users and 30210=3020302*10=3020 total incoming requests, consisting of 1378 appropriate labels and 1642 inappropriate labels.
@@ 5.2. Baselines
5.2.1. Privacy Norms-based Baseline
Privacy norms capture the general population’s accepted appropriateness of information flows. We consider two prompts.
@@ 5.2.1. Privacy Norms-based Baseline
    zero-shot. We construct the prompt to utilize the out-of-the-box privacy norm awareness of LLMs, following from recent prior work on evaluating CI-compliance in LLMs (mireshghallah2024can; shao2024privacylens). The prompt directs the LLM to judge the appropriateness of each information flow based on its knowledge of general privacy norms. The prompts can be found in Appendix A.1.
@@
    privacy norms. We construct the prompt such that the LLM’s privacy judgment is now based on a provided subset of representative privacy norms. These extracted privacy norms are taken directly from the analysis provided in the original papers for each dataset. Table 2 contains a subset of the general privacy norms from both evaluation datasets that were used by the LLMs. The prompt directs the LLM to judge the appropriateness of each information flow based on the provided general privacy norms. The prompts containing the entire set of privacy norms can be found in Appendix A.2.